1) The company Old Co has a current stock price of $210 per share. Last year they paid an
annual dividend of $10 per share. Historically the dividend growth was 5% per year; the next dividend payment will be $10.50. What is the discount rate implied by this firm?
2)A competitor NewCo enters the market (same good, same technology, same market). The
company says they do not plan to pay a dividend for the first five years. Instead, they plan to
use the cash to make strategic acquisitions into new businesses. The company says they will
pay a dividend at the end of year 6 of $20 and then grow the dividend by 5% a year. What is
the present value of NewCo using the discount rate of OldCo?
1
Expert's answer
2020-04-26T18:55:57-0400

1.Gordon Growth Model:

"P=\frac{D}{r-g}"

"210=\\frac{10.5}{r-0.05}\n\n\u200b"

r=0.1, 10%

2.Gordon Growth Model:

"P=\frac{D}{r-g}"

"P=\frac{20}{0.1-0.05}=400"